### v1

@@ Abstract
Adapting pre-trained deep learning models to new and unknown environments is a difficult challenge in underwater acoustic localization. We show that although pre-trained models have performance that suffers from mismatch between the training and test data, they generally exhibit a higher “implied uncertainty” in environments where there is more mismatch. Leveraging this notion of implied uncertainty, we partition the test samples into more certain and less certain sets, and implement an estimation method using the certain samples to improve the labeling for uncertain samples, which helps to adapt the model. We use an efficient method to quantify model prediction uncertainty, and an innovative approach to adapt a pre-trained model to unseen underwater environments at test time. This eliminates the need for labeled data from the target environment or the original training data. This adaptation is enhanced by integrating an independent estimate based on the received signal energy. We validate the approach extensively using real experimental data, as well as synthetic data consisting of model-generated signals with real ocean noise. The results demonstrate significant improvements in model prediction accuracy, underscoring the potential of the method to enhance underwater acoustic localization in diverse, noisy, and unknown environments.
Index Terms:
underwater acoustics, localization, domain adaptation, source hypothesis transfer, mismatch
@@ I Introduction
Deep learning (DL)-based underwater acoustic (UWA) localization algorithms [1, 2, 3, 4] tend to generalize poorly to mismatched environments[3, 5, 6]. To overcome small mismatches between the testing and training sets, domain adaptation (DA) methods match some of the statistics of the testing data to those of the training data[7], which, at inference time, requires access to the training data acquired with known or controlled sources. However, for low-power underwater devices working in a decentralized manner [8, 9, 10], training data will not always be available at the inference device, either due to the communication cost of transferring large datasets between such devices, or security or privacy issues. Therefore, this paper seeks to improve the generalization performance of DL-based models for UWA localization via test-time adaptation (TTA)  [^1] Note that in machine learning parlance, test-time adaptation is also called source-free adaptation, where “source” refers to the training domain. However, to avoid confusion with the acoustic source, we use the term test-time adaptation. [11].
@@
The environmental mismatch problem is known in the machine learning literature as domain shift [14], and has been tackled with approaches including domain adaptation or generalization [14, 11], meta learning [15], transfer learning [16, 17], and data augmentation [18]. Transfer learning requires labels for some of the test data, which are usually not available in UWA problems and encourages pursuit of DA methods. Moreover, although DA and transfer learning have been studied for some cases in underwater acoustic localization [16, 19], [5, 20], more extensive investigations into DA are warranted.
@@
There is a rich literature on TTA including source hypothesis transfer (SHOT) [11], entropy minimization [22], universal TTA [23], contrastive TTA[24], adaptive adversarial network-based TTA [25], and distribution estimation [26]. To perform TTA, we leverage an implied model uncertainty [27, 28, 29, 30]. Although conformal prediction [31] has recently been adopted for uncertainty quantification in data-driven UWA localization [32], it requires a labeled dataset from the test environment for calibration. To explore the effects of mismatch on model uncertainty without any labeled dataset available at the test time, we adopt model uncertainty using mutual information (MUMI) [29]. To avoid the computational burden of MUMI during adaptation, in Section III-B, we cast the ranging problem as a classification task that can leverage computationally simpler implied uncertainty quantification methods such as peakwise uncertainty (PU). Our observations in Section III demonstrate that both MUMI and PU increase with an increase in the amount of mismatch.
@@
The rest of the paper is organized as follows. Section II provides the necessary background on the MFP and CNN approaches for localization. Section III investigates uncertainty quantification in the UWA localization problem. Section IV provides the SHOT approach to adaptation and our proposed method for JSEA. Finally, Section V demonstrates the efficacy of the proposed method on several synthetic and real datasets, followed by some concluding remarks in Section VI.
@@ II Localization Using Sample Covariance Matrices
Notation: Bold letters denote matrices and vectors, $H(\mathbf{y})$ indicates the discrete entropy of a probability mass function (PMF) $\mathbf{y}$, $D_{KL}$ denotes the KL-divergence, $\text{Re}\left(x\right)$ and $\text{Im}\left(x\right)$ respectively denote the real and imaginary parts of $x$, $\mathbf{A}^{\mathsf{T}}$ and $\mathbf{A}^{\mathsf{H}}$ denote respectively the transpose and conjugate transpose of the matrix or vector $\mathbf{A}$. For a set of test points $\mathcal{S}$, the complement is shown by $\mathcal{S}^{\mathsf{c}}=\mathcal{D}_{test}-\mathcal{S}$ and the cardinality is shown by $|\mathcal{S}|$.
@@
where the scalar $r^{(p)}_{l}(f)$ is the complex Fourier coefficient of the $p$-th segment of the received signal on the $l$-th hydrophone at the frequency $f$. For the synthetic data in the simulations, we use the KRAKEN-generated values for $\tilde{r}^{(p)}_{l}(f)$, while for the real data, we take the Fourier transform of the recorded temporal waveform, as explained in Section V-B.
@@ II-A MFP approach
Given that the acoustic source propagates a narrowband signal at a frequency $f$, we use the Bartlett processor [37] as an MFP baseline, which is defined as
@@
where $\tilde{\mathbf{C}}$ is the normalized SCM of the measured data, $\mathbf{\tilde{r}}(f,d)$ is the replica field at frequency $f$ generated by a source at range $d$. For each $d$, the corresponding $\mathbf{\tilde{r}}(f,d)$ (of the CNN approach) from the training set is used at inference time.
@@ II-B CNN approach
The classifier CNN architecture is detailed as follows. As depicted in Fig. 1, the feature extractor part of the model consists of a CNN that takes the real and imaginary parts of the SCM, $\mathbf{x}_{i}=\left[\text{Re}\left(\tilde{\mathbf{C}_{i}}\right),\text{Im}\left(\tilde{\mathbf{C}_{i}}\right)\right]$, as a $2$-channel input (a real tensor of dimension $2\times L\times L$) and a linear layer that generates the feature vector $\boldsymbol{\phi}_{i}$. These features are then fed into a linear classifier to generate the output PMF $\hat{\mathbf{y}}_{i}$. Denoting the region of interest for the source range by $[D_{\min},\;D_{\max}]$, for each training sample $i$, we quantize the range $d_{i}$ by $d^{q}_{i}=\lfloor\frac{d_{i}-D_{\min}}{B}+0.5\rfloor$ to obtain classification outputs with a resolution of $B=100$ m (and consequently, a PMF with bin size $B=100$ m). A one-hot representation of $d^{q}_{i}$ is used in classification tasks. However, since we use mean absolute error (MAE) to evaluate our results, we define the soft label $\mathbf{y}_{i}=[y_{i1},y_{i2},...,y_{iM}]=\eta(d^{q}_{i})$ as
@@
where $M$ is the number of classes and $\sigma$ is a hyperparameter determining the spread of the peak in the soft label. A small $\sigma$ effectively renders the labels one-hot coded while a large $\sigma$ decreases the accuracy by encouraging high-entropy output labels. We empirically conclude that $1\leq\sigma\leq 10$ is appropriate for quantization bins of size $B=100$ m. Figure 2 shows an example of a soft label with $\sigma=5$.
@@
Therefore, the training dataset is denoted by $\mathcal{D}_{tr}=\{(\mathbf{x}_{i},\mathbf{y}_{i})\}_{i=1}^{N_{tr}}$ and the test set that does not include labels is denoted by $\mathcal{D}_{test}=\{\mathbf{x}_{i}\}_{i=1}^{N_{test}}$. We use the loss function $\mathcal{L}_{tr}=\frac{1}{N_{tr}}\sum_{i=1}^{N_{tr}}\mathsf{CE}(\mathbf{y}_{i},\hat{\mathbf{y}}_{i})$ for training, where $\hat{\mathbf{y}}_{i}$ is the network output and $\mathsf{CE}$ is the cross-entropy loss defined by
@@
The CNN structure used herein is similar to that presented in [3] with a $2$-channel input and $3$ convolutional layers with $6$, $38$, and $40$ channels and kernels of sizes $3$, $5$, and $5$, respectively. The output of the feature extraction is a $256$-dimensional vector $\boldsymbol{\phi}_{i}$, which is then fed into the classifier to generate $\hat{\mathbf{y}}_{i}$.
@@ III-A Model Uncertainty using Mutual Information (MUMI)
There are two different sources of uncertainties in data-driven models [29, 30]: 1) epistemic that corresponds to the lack of training data and 2) aleatoric that corresponds to noisy or inherently ambiguous data. To alleviate the effects of mismatch between the training and test data, we focus on the epistemic uncertainty. Consider a dropout network with the weights vector $\mathbf{w}$. Treating $\mathbf{w}$ as a random variable, the network output for a given test sample $\mathbf{x}_{\mathsf{test}}=\left[\text{Re}\left(\tilde{\mathbf{C}}_{\mathsf{test}}\right),\text{Im}\left(\tilde{\mathbf{C}}_{\mathsf{test}}\right)\right]$ is also a random variable. If each realization of $\mathbf{w}$ yields a different output for the same input $\mathbf{x}_{\mathsf{test}}$, the model is considered uncertain about the output. MUMI [29] quantifies the amount of information that we would obtain had we received the true label (acoustic source range) $d_{\mathsf{test}}$ for the corresponding $\mathbf{x}_{\mathsf{test}}$. This is represented in terms of the mutual information [39] $I(d_{\mathsf{test}};\mathbf{w}|\mathcal{D}_{tr},\mathbf{x}_{\mathsf{test}})$, where the conditioning over $\mathcal{D}_{tr}$ is because we have used the training data to obtain $\mathbf{w}$. Expanding this mutual information yields
$I(d_{\mathsf{test}};\mathbf{w}|\mathcal{D}_{tr},\mathbf{x}_{\mathsf{test}})$ $=H\left(p(d_{\mathsf{test}}|\mathcal{D}_{tr},\mathbf{x}_{\mathsf{test}})\right)$
$-\mathbb{E}_{p(\mathbf{w}|\mathcal{D}_{tr})}\left[H\left(p(d_{\mathsf{test}}|\mathcal{D}_{tr},\mathbf{w},\mathbf{x}_{\mathsf{test}})\right)\right],$ (6)
where for a given function $h$, $\mathbb{E}_{p(\mathbf{w}|\mathcal{D}_{tr})}[h(\mathbf{w})]$ denotes the expectation of the random variable $h(\mathbf{w})$ when $\mathbf{w}$ is distributed according to the probability density function $p(\mathbf{w}|\mathcal{D}_{tr})$. Although computing $\mathbb{E}_{p(\mathbf{w}|\mathcal{D}_{tr})}[h(\mathbf{w})]$ is intractable in general, a dropout approximating distribution $q_{\theta}(\mathbf{w}|\mathcal{D}_{tr})$ can efficiently approximate $p(\mathbf{w}|\mathcal{D}_{tr})$ [29], where $\theta$ represents the parameters of a dropout neural network and each realization of the dropout mask corresponds to a sampling from $q_{\theta}(\mathbf{w}|\mathcal{D}_{tr})$ that yields a different weight vector $\mathbf{w}_{j}$. Accordingly, the label likelihood $p(d_{\mathsf{test}}|\mathcal{D}_{tr},\mathbf{x}_{\mathsf{test}})$ in (III-A) can be computed by marginalization over $\mathbf{w}$ and a Monte-Carlo approximation as
[…]
$\triangleq\frac{1}{J}\sum_{j=1}^{J}\hat{p}(\hat{y}_{j}|\mathbf{w}_{j},\mathcal{D}_{tr}\mathbf{x}_{\mathsf{test}})$
@@
where $\mathbf{w}_{j}\sim q_{\theta}(\mathbf{w}|\mathcal{D}_{tr}),j=1,...,J$ are weights sampled from the dropout network with $J$ different dropout realizations, $\hat{y}_{j}$ is the network output corresponding to the weight vector $\mathbf{w}_{j}$, and $\hat{p}$ is the empirical likelihood of the model output. Since the training data has been used to obtain $\mathbf{w}$, there is a Markov chain $\mathcal{D}_{tr}-\mathbf{w}-\hat{y}_{j}$, hence, $\hat{p}(\hat{y}_{j}|\mathcal{D}_{tr},\mathbf{w}_{j},\mathbf{x}_{\mathsf{test}})=\hat{p}(\hat{y}_{j}|\mathbf{w}_{j},\mathbf{x}_{\mathsf{test}})$. Similarly, $p(d_{\mathsf{test}}|\mathbf{w},\mathcal{D}_{tr},\mathbf{x}_{\mathsf{test}})$ can be approximated with $\hat{p}(\hat{y}_{j}|\mathbf{w}_{j},\mathbf{x}_{\mathsf{test}})$, which results in
$\mathbb{E}_{p(\mathbf{w}|\mathcal{D}_{tr})}$ $\left[H\left(p(d_{\mathsf{test}}|\mathcal{D}_{tr},\mathbf{w},\mathbf{x}_{\mathsf{test}})\right)\right]$
@@
$I(d_{\mathsf{test}};\mathbf{w}|\mathcal{D}_{tr},\mathbf{x}_{\mathsf{test}})$ $\approx H\left(\frac{1}{J}\sum_{j=1}^{J}\hat{p}(\hat{y}_{j}|\mathbf{w}_{j},\mathbf{x}_{\mathsf{test}})\right)$
@@
$I(d_{\mathsf{test}};\mathbf{w}|\mathcal{D}_{tr},\mathbf{x}_{\mathsf{test}})\approx H\left(\frac{1}{J}\sum_{j=1}^{J}\hat{p}(\hat{y}_{j}|\mathbf{w}_{j},\mathbf{x}_{\mathsf{test}})\right).$ (9)
Figure 3 shows that when presented with samples from a mismatched environment, a pre-trained regression model will have a higher MUMI. This shows that the average MUMI increases with the increase in depth mismatch or SSP mismatch. This behavior is similar to the increased MUMI of DL models during adversarial attacks discussed in [29]. Also, MUMI is seen to be a monotonically increasing function of the mismatch amount.
@@
(c) Average uncertainty under SSP mismatch.
(d) Average uncertainty under depth mismatch.
Figure 3: MUMI uncertainty of a regression dropout network reveals the mismatch between the training and test environments. These are using the SWellEx-96-inspired synthetic data, consisting of a shallow source with a monotone signal at $f=109$ Hz and all of the results are averaged over $100$ independent realizations of an additive white Gaussian noise (AWGN) channel. Top row shows that the uncertainty is different for each sample and the bottom row shows that, on average, both the uncertainty and the root-mean-squared-error (RMSE) increase with the amount of mismatch.

### v2

@@ Abstract
Adapting pre-trained deep learning models to new and unknown environments remains a major challenge in underwater acoustic localization. We show that although the performance of pre-trained models suffers from mismatch between the training and test data, they generally exhibit a higher uncertainty in environments where there is more mismatch. Additionally, in the presence of environmental mismatch, spurious peaks can appear in the output of classification-based localization approaches, which inspires us to define and use a method to quantify the “implied uncertainty” based on the number of model output peaks. Leveraging this notion of implied uncertainty, we partition the test samples into sets with more certain and less certain samples, and implement a method to adapt the model to new environments by using the certain samples to improve the labeling for uncertain samples. Thus, using this efficient method for model uncertainty quantification, we showcase an innovative approach to adapt a pre-trained model to unseen underwater environments at test time. This eliminates the need for labeled data from the target environment or the original training data. This adaptation is enhanced by integrating an independent estimate based on the received signal energy. We validate the approach extensively using real experimental data, as well as synthetic data consisting of model-generated signals with real ocean noise. The results demonstrate significant improvements in model prediction accuracy, underscoring the potential of the method to enhance underwater acoustic localization in diverse, noisy, and unknown environments.
@@ I Introduction
Underwater acoustic (UWA) source localization refers to determining all or some of the coordinates of an underwater acoustic source using the corresponding recorded signals [1, 2, 3, 4, 5, 6]. Source ranging refers to estimating the distance of a source and can be achieved using conventional methods such as matched field processing (MFP) [1, 2] or deep learning (DL)-based methods [3, 7, 8]. DL-based UWA localization algorithms [3, 9, 7, 10] tend to generalize poorly to mismatched environments[7, 11, 12]. To overcome small mismatches between the testing and training sets, domain adaptation (DA) methods have been proposed that match some of the statistics of the testing data to those of the training data[13], which, at inference time, requires access to the training data acquired with known or controlled sources. However, for low-power or low storage-capacity underwater devices working in a decentralized manner [14, 15, 16], training data will not always be available at the inference device, either due to the communication cost of transferring large datasets between such devices, or security or privacy issues. Therefore, this paper seeks to improve the generalization performance of DL-based models for UWA localization via test-time adaptation (TTA)  [^1] Note that in machine learning parlance, test-time adaptation is also called source-free adaptation, where “source” refers to the training domain. However, to avoid confusion with the acoustic source, we use the term test-time adaptation. [17].
@@
The environmental mismatch problem is known in the machine learning literature as domain shift [19], and has been tackled with approaches including DA or generalization [19, 17], meta learning [20], transfer learning [21, 22], and data augmentation [23]. Transfer learning requires labels for some of the test data, which are usually not available in UWA problems and encourages pursuit of DA methods. Moreover, although DA and transfer learning have been studied for some cases in underwater acoustic localization [21, 24], [11, 25], more extensive investigations into DA are warranted.
@@
There is a rich literature on TTA including source hypothesis transfer (SHOT) [17], entropy minimization [27], universal TTA [28], contrastive TTA[29], adaptive adversarial network-based TTA [30], and distribution estimation [31]. To perform TTA, we leverage an implied model uncertainty [32, 33, 34, 35]. Although conformal prediction [36] has recently been adopted for uncertainty quantification in data-driven UWA localization [37], it requires a labeled dataset from the test environment for calibration. To explore the effects of mismatch on model uncertainty without any labeled dataset available at the test time, we adopt model uncertainty using mutual information (MUMI) [34] as a metric. To avoid the computational burden of MUMI during adaptation, in Section III-B, we cast the ranging problem as a classification task that can leverage computationally simpler implied uncertainty quantification methods such as peakwise uncertainty (PU). Our observations in Section III demonstrate that both MUMI and PU increase with an increase in the amount of mismatch, validating their utility.
@@
The rest of the paper is organized as follows. Section II provides the necessary background on the MFP and CNN approaches for localization. Section III investigates uncertainty quantification in the UWA localization problem. Section IV provides the SHOT approach to adaptation and our proposed method for JSEA. Finally, Section V demonstrates the efficacy of the proposed method on several synthetic and real datasets, followed by discussions on potential extensions and limitations of JSEA in Section VI and some concluding remarks in Section VII.
@@ II Localization Using Sample Covariance Matrices
Notation: Bold letters denote matrices and vectors, $H(\mathbf{y})$ indicates the discrete entropy of a probability mass function (PMF) $\mathbf{y}$, $D_{KL}$ denotes the KL-divergence, $\real{x}$ and $\text{Im}\left(x\right)$ respectively denote the real and imaginary parts of $x$, Superscripts T and H denote respectively the transpose and conjugate transpose of the matrices or vectors. For a set of test points $\mathcal{S}$, the complement is represented by $\mathcal{S}^{\mathsf{c}}=\mathcal{D}_{\text{test}}-\mathcal{S}$ and the cardinality is shown by $|\mathcal{S}|$.
@@
where the scalar $r^{(p)}_{l}(f)$ is the complex Fourier coefficient of the $p$-th segment of the received signal on the $l$-th hydrophone at the frequency $f$. For the synthetic data in the simulations, we use the signals generated by the KRAKEN normal-mode model[41] for $\tilde{r}^{(p)}_{l}(f)$, while for the real data, we take the Fourier transform of the recorded temporal waveform and use the Fourier coefficients at the frequency of interest, as explained in Section V-B.
@@ II-A MFP approach
Given that the acoustic source emits a narrowband signal at a frequency $f$, we use the Bartlett processor [42] as an MFP baseline, which is defined as
@@
where $\tilde{\mathbf{C}}$ is the normalized SCM of the measured data, and $\mathbf{\tilde{r}}(f,d)$ is the replica field at frequency $f$ generated by a source at range $d$. For each $d$, the corresponding $\mathbf{\tilde{r}}(f,d)$ (of the CNN approach) from the training set is used at inference time.
@@ II-B CNN approach
[…]
@@
where $M$ is the number of classes and $\sigma$ is a hyperparameter determining the spread of the peak in the soft label. A small $\sigma$ effectively renders the labels one-hot coded while a large $\sigma$ decreases the accuracy by encouraging high-entropy output labels. We empirically conclude that $1\leq\sigma\leq 10$ is appropriate for quantization bins of size $B=100$ m. Fig. 2 shows an example of a soft label with $\sigma=5$.
@@
Therefore, the training dataset is denoted by $\mathcal{D}_{\text{tr}}=\{(\mathbf{x}_{i},\mathbf{y}_{i})\}_{i=1}^{N_{\text{tr}}}$ and the test set that does not include labels is denoted by $\mathcal{D}_{\text{test}}=\{\mathbf{x}_{i}\}_{i=1}^{N_{\text{test}}}$. We use the loss function $\mathcal{L}_{tr}=\frac{1}{N_{\text{tr}}}\sum_{i=1}^{N_{\text{tr}}}\mathsf{CE}(\mathbf{y}_{i},\hat{\mathbf{y}}_{i})$ for training, where $\hat{\mathbf{y}}_{i}$ is the network output and $\mathsf{CE}$ is the cross-entropy loss defined by
@@
The CNN structure used herein is similar to that presented in [7] with a $2$-channel input and $3$ convolutional layers with $6$, $38$, and $40$ channels and kernels of sizes $3$, 55, and $5$, respectively. The output of the feature extraction is a $256$-dimensional vector $\boldsymbol{\phi}_{i}$, which is then fed into the classifier to generate $\hat{\mathbf{y}}_{i}$.
@@ III-A Model Uncertainty using Mutual Information (MUMI)
There are two different sources of uncertainties in data-driven models [34, 35]: 1) epistemic that corresponds to the lack of training data and 2) aleatoric that corresponds to noisy or inherently ambiguous data. To alleviate the effects of mismatch between the training and test data, we focus on the epistemic uncertainty. Consider a dropout network with the weights vector $\mathbf{w}$. Treating $\mathbf{w}$ as a random variable, the network output for a given test sample $\mathbf{x}_{\mathsf{test}}=\left[\real{\tilde{\mathbf{C}}_{\mathsf{test}}},\text{Im}\left(\tilde{\mathbf{C}}_{\mathsf{test}}\right)\right]$ is also a random variable. If each realization of $\mathbf{w}$ yields a different output for the same input $\mathbf{x}_{\mathsf{test}}$, the model is considered uncertain about the output. MUMI [34] quantifies the amount of information that we would obtain had we received the true label (acoustic source range) $d_{\mathsf{test}}$ for the corresponding $\mathbf{x}_{\mathsf{test}}$. This is represented in terms of the mutual information [44] $I(d_{\mathsf{test}};\mathbf{w}|\mathcal{D}_{\text{tr}},\mathbf{x}_{\mathsf{test}})$, where the conditioning over $\mathcal{D}_{\text{tr}}$ is because we have used the training data to obtain $\mathbf{w}$. Expanding this mutual information yields
$I(d_{\mathsf{test}};\mathbf{w}|\mathcal{D}_{\text{tr}},\mathbf{x}_{\mathsf{test}})$ $=H\left(p(d_{\mathsf{test}}|\mathcal{D}_{\text{tr}},\mathbf{x}_{\mathsf{test}})\right)$
$-\mathbb{E}_{p(\mathbf{w}|\mathcal{D}_{\text{tr}})}\left[H\left(p(d_{\mathsf{test}}|\mathcal{D}_{\text{tr}},\mathbf{w},\mathbf{x}_{\mathsf{test}})\right)\right],$ (6)
where for a given function $h$, $\mathbb{E}_{p(\mathbf{w}|\mathcal{D}_{\text{tr}})}[h(\mathbf{w})]$ denotes the expectation of the random variable $h(\mathbf{w})$ when $\mathbf{w}$ is distributed according to the probability density function $p(\mathbf{w}|\mathcal{D}_{\text{tr}})$. Although computing $\mathbb{E}_{p(\mathbf{w}|\mathcal{D}_{\text{tr}})}[h(\mathbf{w})]$ is intractable in general, a dropout approximating distribution $q_{\theta}(\mathbf{w}|\mathcal{D}_{\text{tr}})$ can efficiently approximate $p(\mathbf{w}|\mathcal{D}_{\text{tr}})$ [34], where $\theta$ represents the parameters of a dropout neural network and each realization of the dropout mask corresponds to a sampling from $q_{\theta}(\mathbf{w}|\mathcal{D}_{\text{tr}})$ that yields a different weight vector $\mathbf{w}_{j}$. Accordingly, the label likelihood $p(d_{\mathsf{test}}|\mathcal{D}_{\text{tr}},\mathbf{x}_{\mathsf{test}})$ in (III-A) can be computed by marginalization over $\mathbf{w}$ and a Monte-Carlo approximation as
$p(d_{\mathsf{test}}|\mathcal{D}_{\text{tr}},\mathbf{x}_{\mathsf{test}})$ $\approx\int p(d_{\mathsf{test}}|\mathbf{w},\mathcal{D}_{\text{tr}},\mathbf{x}_{\mathsf{test}})q_{\theta}(\mathbf{w}|\mathcal{D}_{\text{tr}})d\mathbf{w}$
$\triangleq\frac{1}{J}\sum_{j=1}^{J}\hat{p}(\hat{y}_{j}|\mathbf{w}_{j},\mathcal{D}_{\text{tr}},\mathbf{x}_{\mathsf{test}})$
@@
where $\mathbf{w}_{j}\sim q_{\theta}(\mathbf{w}|\mathcal{D}_{\text{tr}}),j=1,...,J$ are weights sampled from the dropout network with $J$ different dropout realizations, $\hat{y}_{j}$ is the network output corresponding to the weight vector $\mathbf{w}_{j}$, and $\hat{p}$ is the empirical likelihood of the model output. Since the training data has been used to obtain $\mathbf{w}$, there is a Markov chain $\mathcal{D}_{\text{tr}}-\mathbf{w}-\hat{y}_{j}$, hence, $\hat{p}(\hat{y}_{j}|\mathcal{D}_{\text{tr}},\mathbf{w}_{j},\mathbf{x}_{\mathsf{test}})=\hat{p}(\hat{y}_{j}|\mathbf{w}_{j},\mathbf{x}_{\mathsf{test}})$. Similarly, $p(d_{\mathsf{test}}|\mathbf{w},\mathcal{D}_{\text{tr}},\mathbf{x}_{\mathsf{test}})$ can be approximated with $\hat{p}(\hat{y}_{j}|\mathbf{w}_{j},\mathbf{x}_{\mathsf{test}})$, which results in
$\mathbb{E}_{p(\mathbf{w}|\mathcal{D}_{\text{tr}})}$ $\left[H\left(p(d_{\mathsf{test}}|\mathcal{D}_{\text{tr}},\mathbf{w},\mathbf{x}_{\mathsf{test}})\right)\right]$
@@
$I(d_{\mathsf{test}};\mathbf{w}|\mathcal{D}_{\text{tr}},\mathbf{x}_{\mathsf{test}})$ $\approx H\left(\frac{1}{J}\sum_{j=1}^{J}\hat{p}(\hat{y}_{j}|\mathbf{w}_{j},\mathbf{x}_{\mathsf{test}})\right)$
@@
$I(d_{\mathsf{test}};\mathbf{w}|\mathcal{D}_{\text{tr}},\mathbf{x}_{\mathsf{test}})\approx H\left(\frac{1}{J}\sum_{j=1}^{J}\hat{p}(\hat{y}_{j}|\mathbf{w}_{j},\mathbf{x}_{\mathsf{test}})\right).$ (9)
Fig. 3 shows that when presented with samples from a mismatched environment, a pre-trained regression model will have a higher MUMI. This shows that the average MUMI increases with the increase in depth mismatch or SSP mismatch. This behavior is similar to the increased MUMI of DL models during adversarial attacks discussed in [34]. Also, MUMI is seen to be a monotonically increasing function of the amount of mismatch, validating its use as a metric of uncertainty.
@@
(c) Average uncertainty over all ranges considered under SSP mismatch.
(d) Average uncertainty over all ranges considered under depth mismatch.
Figure 3: MUMI uncertainty of a regression dropout network reveals the mismatch between the training and test environments. These are using the SWellEx-96-inspired synthetic data, consisting of a shallow source with a monotone signal at $f=109$ Hz and all of the results are averaged over $100$ independent realizations of an additive white Gaussian noise (AWGN) channel. Top row shows that the uncertainty is different for each source range and the bottom row shows that, on average, both the uncertainty and the root-mean-squared-error (RMSE) increase with the amount of mismatch.